𝐠+2is afactor of the polynomial 𝐵3 − 2𝐲 + 𝐵− 18.When this polynomialis divided by 𝐠+ 1the remainder is 18.a. Find the values of 𝐠and 𝐮b. With thesevalues find the remaining factorsof the polynomial.c. Hence,solve the equation 𝐵3 −2𝐲 +𝐵 = 18.d. Sketch the graph of the polynomialusing the values of 𝐠and𝑹ou have found.


1
Expert's answer
2021-03-09T07:20:36-0500

The given Polynomial is "B^3-2y+B-18."

As g+2 is a factor of polynomial


Then (-2) is a factor od polynomial


"\Rightarrow (-2)^3-2y+(-2)-18=0\\\Rightarrow 2y=-8-2-18\\\Rightarrow y=-14"


Put this value in Polyomial

we get polynomail as "B^3+28+B-18"


i.e. "B^3+B+10=0"